# Dynamic Pre-training: Towards Efficient and Scalable All-in-One Image Restoration

### v1

@@ Abstract
All-in-one image restoration tackles different types of degradations with a unified model instead of having task-specific, non-generic models for each degradation. The requirement to tackle multiple degradations using the same model can lead to high-complexity designs with fixed configuration that lack the adaptability to more efficient alternatives. We propose DyNet, a dynamic family of networks designed in an encoder-decoder style for all-in-one image restoration tasks. Our DyNet can seamlessly switch between its bulkier and lightweight variants, thereby offering flexibility for efficient model deployment with a single round of training. This seamless switching is enabled by our weights-sharing mechanism, forming the core of our architecture and facilitating the reuse of initialized module weights. Further, to establish robust weights initialization, we introduce a dynamic pre-training strategy that trains variants of the proposed DyNet concurrently, thereby achieving a 50% reduction in GPU hours. To tackle the unavailability of large-scale dataset required in pre-training, we curate a high-quality, high-resolution image dataset named Million-IRD having 2M image samples. We validate our DyNet for image denoising, deraining, and dehazing in all-in-one setting, achieving state-of-the-art results with 31.34% reduction in GFlops and a 56.75% reduction in parameters compared to baseline models. The source codes and trained models are available at https://github.com/akshaydudhane16/DyNet.
Keywords:
All-in-one Image Restoration Foundation model
@@ 1 Introduction
The image restoration (IR) task seeks to improve low-quality input images. Despite several advancements in IR, diverse degradation types and severity levels present in images continue to pose a significant challenge. The majority of existing methods [46, 31, 13, 8, 15, 52, 3, 14, 47] learn image priors implicitly, requiring separate network training for different degradation types, levels, and datasets. Further, these methods require a prior knowledge of image degradation for effective model selection during testing, thereby lacking generality to cater diverse degradations.
@@
All-in-one restoration aims to restore images with an unknown degradation using a single unified model. Recent advancements, such as AirNet [26] and PromptIR [36], have addressed the all-in-one restoration challenge by employing contrastive learning and implicit visual prompting techniques, respectively. Specifically, the state-of-the-art PromptIR [36] employs an implicit prompting to learn degradation-aware prompts on the decoder side, aiming to refine the decoder features. This approach, while intriguing, does not extend to the refinement of encoder features, which are left unprocessed. Despite the interesting idea of implicit prompting, it poses a considerable challenge for deployment due to its poor computational efficiency with 37M parameters and 243 GFlops to process a single 224$\times$224 sized image. The practical application of such models is therefore challenging due to the high computational requirements, especially in resource-constraint hand-held devices. However, choosing lightweight models invariably leads to a trade-off between accuracy and efficiency [56, 28]. Most of these models aim to reduce parameter counts by strided convolutions and feature channel splitting [23, 28, 22]. Some strategies also operate within the frequency domain to reduce the computational demands linked to attention mechanisms [62] or non-local operations [18] while few methods like [9, 49] partition the feature space for efficient processing, or split the attention across dimensions to improve the computational efficiency [61, 9].
@@
Towards optimizing all-in-one IR efficiency without sacrificing performance, in this paper, we introduce a novel weights-sharing mechanism. In this scheme, the weights of a network module are shared with its subsequent modules in a series. This approach substantially reduces the number of parameters, resulting in a more streamlined network architecture. We deploy the proposed weights-sharing mechanism in an encoder-decoder style architecture named Dynamic Network (DyNet). Apart from the computational efficiency, the proposed DyNet provides exceptional flexibility as by merely changing the module weights reuse frequency, one can easily adjust the network depth and correspondingly switch between its bulkier and light-weight variants during training.
@@
Figure 1: Left: At any given encoder-decoder level, (a) Series of transformer blocks in existing PromptIR [36], (b) The proposed transformer block weights sharing mechanism in our DyNet-L. We initialize one transformer block at each encoder-decoder level and shared its weights with the subsequent $N$ transformer blocks. (c) Our lightweight variant DyNet-S. Here, the number of transformer blocks are halved and weights of the first transformer block are shared with $N/2$ subsequent transformer blocks. Right: A plot of average PSNR in All-in-one restoration setting vs GFlops and parameters (in millions). The proposed DyNet-S achieves boost of 0.59 dB with 31.34% reduction in GFlops and a 56.75% reduction in parameters compared to the baseline PromptIR [36].
@@
Through the synergy of these techniques, our proposed DyNet achieves an average gain of 0.82 dB for image de-noising, deraining, and dehazing within an all-in-one setting, with 31.34% reduction in GFlops and a 56.75% reduction in network parameters compared to the baseline; refer Fig. 1.
@@ 2 Related Work
Image restoration seeks to restore images from their degraded low-quality versions, with the restoration process varying considerably across different tasks. Research in image restoration has predominantly concentrated on addressing single degradation challenges [52, 40, 11, 53, 38, 58, 45, 34, 60]. Although significant progress has been made in single degradation restoration techniques, research in multi-degradation restoration using a unified model is still relatively under-explored. Multi-degradation restoration, however, provides more practical advantages as the information about input degradation type is not required to select the task-specific model and it also enhances model storage efficiency over single-degradation restoration.
@@
The critical challenge in multi-task image restoration is developing a single model capable of addressing various types of degradation and precisely restoring low-quality images. The IPT model [7], for instance, relies on prior knowledge of the corruption affecting the input image, utilizing a pre-trained transformer backbone equipped with distinct encoders and decoder heads tailored for multi-task restoration. However, in blind image restoration, we do not have such prior information about the degradation. In this direction, Li et al. [26] propose a unified network for multiple tasks such as denoising, deraining, and dehazing, employing an image encoder refined through contrastive learning. However, it requires a two-stage training process, where the outcome of contrastive learning relies on selecting positive and negative pairs and the availability of large data samples. Further, motivated by easy implementation and broad applicability of prompts, PromptIR [36] introduces a concept where degradation-aware prompts are learned implicitly within an encoder-decoder architecture for all-in-one image restoration. PromptIR applies implicit prompting to the decoder side to refine the decoder features. However, it does not extend to the refinement of encoder features, which are left unprocessed.
@@
Although the above-discussed methods tackle multi-task image restoration, they suffer from low efficiency due to a large number of parameters and substantial GFlops consumption. Consequently, deploying these models in environments where efficiency is crucial factor becomes difficult due to their demanding computational needs. However, opting for lightweight models introduces an inevitable accuracy-efficiency trade-off [56]. The primary cause is that many current methods focus on reducing parameter counts through techniques like strided convolutions and feature channel splitting [23, 28, 22]. Additionally, few approaches work within the frequency domain to reduce the computational burden associated with attention mechanisms [62] or non-local operations [18]. Furthermore, methods like [9, 49] partition the feature space for efficient processing, while [61, 9] split the attention across dimensions to improve the computational efficiency. In contrast, our proposed approach focuses on an efficient and scalable all-in-one image restoration model that incorporates a module weights-sharing mechanism. We demonstrate that strategic fundamental adjustments to the network architecture result in substantial performance improvements.
@@
Figure 2: The overall pipeline of the proposed Dynamic Network (DyNet) for all-in-one image restoration. DyNet enhances a low-quality input image using a four-level encoder-decoder architecture. A distinctive aspect of DyNet lies in its weight-sharing strategy: at each level, the initial transformer block’s weights are sequentially passed to subsequent blocks, significantly reducing the network’s parameters and enhancing its flexibility. This approach allows for easy modification of DyNet’s complexity, toggling between more bulkier or lightweight variants by adjusting the frequency of weight sharing among the blocks at each encoder-decoder level. Moreover, we maintain the encoder-decoder feature consistency by implicitly learning degradation-aware prompts at skip connections rather than on the decoder side as in PromptIR [36].
@@ 3 Proposed Approach
Existing state-of-the-art all-in-one image restoration (IR) approaches incur a high computational footprint and do not provide the flexibility to support models with variable depths during training. Towards more accurate, but lightweight and flexible architecture design for all-in-one IR, this work proposes a Dynamic Network Architecture (Sec. 3.1) and a Dynamic Pretraining Strategy (Sec. 3.2).
@@ 3.1 Dynamic Network (DyNet) Architecture
Our lightweight architecture is based on a novel weight-sharing mechanism for IR models. This mechanism allows the network module’s weights to be reused across subsequent modules in sequence, thereby significantly reducing the total number of parameters and leading to a more efficient network structure. We implement this weight-sharing approach within an encoder-decoder style architecture, which we term the Dynamic Network (DyNet). Within DyNet, at each encoder-decoder level, module weights are shared across a pre-specified number of subsequent modules. This not only enhances computational efficiency but also grants remarkable flexibility to the architecture. By simply altering the frequency of weight sharing, users can easily modify the network’s depth, seamlessly transitioning between bulkier or lightweight variants of DyNet.
@@
The overall pipeline of the proposed DyNet is shown in Fig. 2. The proposed DyNet begins by extracting low-level features $\mathbf{F_{0}}$in $\mathbb{R}^{H\times W\times C}$ from a given degraded input image $\mathbf{{I}}$in $\mathbb{R}^{H\times W\times 3}$ through a convolution operation. Subsequently, the feature embedding undergoes a 4-level hierarchical encoder-decoder, gradually transforming into deep latent features $\mathbf{F}_{l}$in $\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 8C}$. We utilize the existing Restormer’s [52] Transformer block as a fundamental feature extraction module. For every level of the encoder and decoder, we initialize the weights for a first transformer block, which are subsequently reused across the subsequent blocks in a series up to the specified frequency at that level. For example, at a given encoder/decoder level, $w^{1}$ denotes weights of the first transformer block, we can define the output $\mathbf{F}_{out}$ of that encoder/decoder level, given input features $\mathbf{F}_{in}$ as,
@@
We gradually increase the reuse frequency of the transformer block from the top level to the bottom level, thereby increasing the network depth. The multi-level encoder systematically reduces spatial resolution while increasing channel capacity, facilitating the progressive extraction of low-resolution latent features from a high-resolution input image. Subsequently, the multi-level decoder gradually restores the high-resolution output from these low-resolution latent features. To enhance the decoding process, we integrate prompt blocks [36], which learn degradation-aware prompts implicitly through the prompt generation module (PGM) followed by the prompt-interaction module (PIM). Unlike the existing PromptIR [36], our approach incorporates degradation-aware implicit prompt blocks at skip connections. This implicit prompting via skip connections enables the transfer of degradation-aware encoder features to the decoder side, thereby enhancing the restoration process. This fundamental correction in the network design gives us significant improvement for all-in-one image restoration tasks, as detailed in the experimental section (Sec. 5).
@@
At each encoder-decoder level, by simply changing the module weights reuse frequency ($f$), we obtain DyNet’s bulky and lightweight variants. Fig. 2 shows the core architecture of our DyNet. At each encoder-decoder level, we initialize weights ($w^{1}$) for the first transformer block [52] and reuse them for the subsequent blocks. We vary the reuse frequency at each encoder-decoder level and obtain our bulkier and lightweight variants. DyNet comprises four encoder-decoder levels, with the bulky variant (DyNet-L) employing reuse frequencies of $f=[4,6,6,8]$ at encoder-decoder level 1 to level 4, respectively. In contrast, the lightweight variant (DyNet-S) applies weights reuse frequencies $f=[2,3,3,4]$ for encoder-decoder level 1 to level 4, respectively. Further, we incorporate a prompt block [36] at each skip connection, which enhances the encoder features being transferred to the decoder from the encoder. Overall, there are three prompt blocks deployed at each level, each consisting of five prompt components.
Figure 3: The proposed dynamic pre-training strategy is shown. Given a clean image, we create a degraded version by injecting noise (Gaussian or Random), JPEG artifacts and random masking within the same image. Two variants (small and large) of our DyNet are then trained in parallel branches to reconstruct the clean image from masked degraded inputs. Notably, the weights are shared between both variants (DyNet-S and DyNet-L) since they are based on the same architecture but with an intra-network weight-sharing scheme with varying frequencies of block repetition. One of the two parallel branches is randomly activated in a single forward pass of the model (Flag being the binary indicator variable to show branch activation). The dotted lines show the network path is inactivated (Flag=0). An L1 loss is used to calculate pixel differences between reconstructed outputs and targets to update the shared weights of both branches. Both the inter and intra-model weight-sharing as well as random activation of branches lead to a significant reduction in GPU hours required for pre-training.

### v2

@@ Abstract
All-in-one image restoration tackles different types of degradations with a unified model instead of having task-specific, non-generic models for each degradation. The requirement to tackle multiple degradations using the same model can lead to high-complexity designs with fixed configuration that lack the adaptability to more efficient alternatives. We propose DyNet, a dynamic family of networks designed in an encoder-decoder style for all-in-one image restoration tasks. Our DyNet can seamlessly switch between its bulkier and lightweight variants, thereby offering flexibility for efficient model deployment with a single round of training. This seamless switching is enabled by our weights-sharing mechanism, forming the core of our architecture and facilitating the reuse of initialized module weights. Further, to establish robust weights initialization, we introduce a dynamic pre-training strategy that trains variants of the proposed DyNet concurrently, thereby achieving a 50% reduction in GPU hours. Our dynamic pre-training strategy eliminates the need for maintaining separate checkpoints for each variant, as all models share a common set of checkpoints, varying only in model depth. This efficient strategy significantly reduces storage overhead and enhances adaptability. To tackle the unavailability of large-scale dataset required in pre-training, we curate a high-quality, high-resolution image dataset named Million-IRD, having 2M image samples. We validate our DyNet for image denoising, deraining, and dehazing in all-in-one setting, achieving state-of-the-art results with 31.34% reduction in GFlops and a 56.75% reduction in parameters compared to baseline models. The source codes and trained models are available at https://github.com/akshaydudhane16/DyNet.
@@ 1 Introduction
The image restoration (IR) task seeks to improve low-quality input images. Despite several advancements in IR, diverse degradation types and severity levels present in images continue to pose a significant challenge. The majority of existing methods (Tu et al., 2022; Zamir et al., 2022; Wang et al., 2021; Zamir et al., 2021; 2020b; Chen et al., 2022) learn image priors implicitly, requiring separate network training for different degradation types, levels, and datasets. Further, these methods require a prior knowledge of image degradation for effective model selection during testing, thereby lacking generality to cater diverse degradations.
@@
All-in-one restoration aims to restore images with an unknown degradation using a single unified model. Recent advancements, such as AirNet (Li et al., 2022) and PromptIR (Potlapalli et al., 2023), have addressed the all-in-one restoration challenge by employing contrastive learning and implicit visual prompting techniques, respectively. Specifically, the state-of-the-art PromptIR (Potlapalli et al., 2023) employs an implicit prompting to learn degradation-aware prompts on the decoder side, aiming to refine the decoder features. This approach, while intriguing, does not extend to the refinement of encoder features, which are left unprocessed. Despite the interesting idea of implicit prompting, it poses a considerable challenge for deployment due to its poor computational efficiency with 37M parameters and 243 GFlops to process a single 224$\times$224 sized image. The practical application of such models is therefore challenging due to the high computational requirements, especially in resource-constraint hand-held devices. However, choosing lightweight models invariably leads to a trade-off between accuracy and efficiency (Zhang et al., 2020a; Li et al., 2023b). Most of these models aim to reduce parameter counts by strided convolutions and feature channel splitting (Kong et al., 2022; Li et al., 2023b; Kligvasser et al., 2018). Some strategies also operate within the frequency domain to reduce the computational demands linked to attention mechanisms (Zhou et al., 2023) or non-local operations (Guo et al., 2021) while few methods like (Cui & Knoll, 2023; Xie et al., 2021) partition the feature space for efficient processing, or split the attention across dimensions to improve the computational efficiency (Zhao et al., 2023; Cui & Knoll, 2023).
@@
To optimize all-in-one IR efficiency without sacrificing performance, in this paper, we propose a novel weights-sharing mechanism. In this scheme, the weights of a network module are shared with its subsequent modules in a series. This approach substantially reduces the number of parameters, resulting in a more streamlined network architecture. We deploy the proposed weights-sharing mechanism in an encoder-decoder style architecture named Dynamic Network (DyNet). Apart from the computational efficiency, the proposed DyNet provides exceptional flexibility as by merely changing the module weights reuse frequency; one can easily adjust the network depth and correspondingly switch between its bulkier and lightweight variants. Model variants share the same checkpoints, differing only in depth, eliminating the need for separate checkpoints and saving disk space.
@@
Figure 1: Left: At any given encoder-decoder level: (a) Transformer blocks in PromptIR, (b) and (c) Our DyNet-L and DyNet-S use the proposed weights sharing mechanism, initializing one transformer block and sharing its weights with subsequent blocks. Right: A plot of average PSNR in All-in-one IR setting vs GFlops and parameters (in millions). Our DyNet-S boosts performance by 0.43 dB, while reducing GFlops by 31.34% and parameters by 56.75% compared to PromptIR.
@@
Through the synergy of these techniques, our proposed DyNet achieves an average gain of 0.68 dB for image de-noising, deraining, and dehazing within an all-in-one setting, with 31.34% reduction in GFlops and a 56.75% reduction in network parameters compared to the baseline; refer Fig. 1.
@@ 2 Related Work
Image restoration (IR) seeks to restore images from their degraded low-quality versions, with the restoration process varying considerably across different tasks. Research in IR has predominantly concentrated on addressing single degradation challenges (Zamir et al., 2022; Ren et al., 2020; Dong et al., 2020a; Zamir et al., 2020a; Ren et al., 2021; Zhang et al., 2017b; Tsai et al., 2022; Nah et al., 2022; Zhang et al., 2020b). Although significant progress has been made in single degradation restoration techniques, research in multi-degradation restoration using a unified model is still relatively under-explored. Multi-degradation restoration, however, provides more practical advantages as the information about input degradation type is not required to select the task-specific model and it also enhances model storage efficiency over single-degradation restoration.
@@
The critical challenge in multi-task IR is developing a single model capable of addressing various types of degradation and precisely restoring low-quality images. The IPT model (Chen et al., 2021), for instance, relies on prior knowledge of the corruption affecting the input image, utilizing a pre-trained transformer backbone equipped with distinct encoders and decoder heads tailored for multi-task restoration. However, in blind IR, we do not have such prior information about the degradation. In this direction,  (Li et al., 2022) propose a unified network for multiple tasks such as denoising, deraining, and dehazing, employing an image encoder refined through contrastive learning. However, it requires a two-stage training process, where the outcome of contrastive learning relies on selecting positive and negative pairs and the availability of large data samples. Further, motivated by easy implementation and broad applicability of prompts, PromptIR (Potlapalli et al., 2023) introduces a concept where degradation-aware prompts are learned implicitly within an encoder-decoder architecture for all-in-one IR. PromptIR uses implicit prompting to refine decoder features but does not extend this refinement to the encoder features, leaving them unprocessed.
@@
Although the above-discussed methods tackle multi-task IR, they suffer from low efficiency due to a large number of parameters and substantial GFlops consumption. Consequently, deploying these models in environments where efficiency is crucial factor becomes difficult due to their demanding computational needs. However, opting for lightweight models introduces an inevitable accuracy-efficiency trade-off (Zhang et al., 2020a). The primary cause is that existing methods (Kong et al., 2022; Li et al., 2023b; Kligvasser et al., 2018) focus on reducing parameter counts through techniques like strided convolutions and feature channel splitting. Further, few approaches work in frequency domain to reduce the computational burden associated with attention mechanisms (Zhou et al., 2023) or non-local operations (Guo et al., 2021). Furthermore, methods like (Cui & Knoll, 2023; Xie et al., 2021) partition the feature space for efficient processing, while (Zhao et al., 2023; Cui & Knoll, 2023) split the attention across dimensions to improve the computational efficiency. In contrast, our approach focuses on an efficient and scalable all-in-one image restoration model incorporating a weights-sharing mechanism. We demonstrate that strategic fundamental adjustments to the network architecture result in substantial performance improvements.
@@
Figure 2: The proposed Dynamic Network (DyNet) pipeline for all-in-one image restoration. DyNet enhances a low-quality input image using a 4-level encoder-decoder architecture. A distinctive aspect of DyNet lies in its weight-sharing strategy. At each level, the initial transformer block’s weights are shared with the subsequent blocks, significantly reducing the network parameters and enhancing its flexibility. This approach allows for easy adjustment of DyNet’s complexity, switching between large and small variants by modifying the frequency of weight sharing across the encoder-decoder blocks. Moreover, we maintain the encoder-decoder feature consistency by implicitly learning degradation-aware prompts at skip connections rather than on the decoder side as in PromptIR.
@@ 3 Proposed Approach
Existing SoTA all-in-one image restoration (IR) approaches offer a high computational footprint and do not provide the flexibility to change model complexity ”on-the-go” during deployment. Towards more accurate, but lightweight and flexible architecture design for all-in-one IR, this work proposes a Dynamic Network Architecture (Sec. 3.1) and a Dynamic Pretraining Strategy (Sec. 3.2).
@@ 3.1 Dynamic Network (DyNet) Architecture
Our network architecture is based on a novel weight-sharing mechanism for IR models. This mechanism allows the network module’s weights to be reused across subsequent modules in sequence, thereby significantly reducing the total number of parameters and leading to a more efficient network structure. We implement this weight-sharing approach within an encoder-decoder style architecture, which we term the Dynamic Network (DyNet). Within DyNet, at each encoder-decoder level, module weights are shared across a pre-specified number of subsequent modules. This not only enhances computational efficiency but also grants remarkable flexibility to the architecture. By simply altering the frequency of weight sharing, users can easily modify the network’s depth, seamlessly transitioning between bulkier or lightweight variants of DyNet.
@@
The proposed DyNet pipeline is shown in Fig. 2. The proposed DyNet begins by extracting low-level features $\mathbf{F_{0}}$in $\mathbb{R}^{H\times W\times C}$ from a given degraded input image $\mathbf{{I}}$in $\mathbb{R}^{H\times W\times 3}$ through a convolution operation. Subsequently, the feature embedding undergoes a 4-level hierarchical encoder-decoder, gradually transforming into deep latent features $\mathbf{F}_{l}$in $\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 8C}$. We utilize the existing Restormer’s (Zamir et al., 2022) Transformer block as a fundamental feature extraction module. For every level of the encoder and decoder, we initialize the weights for a first transformer block, which are subsequently reused across the subsequent blocks in a series up to the specified frequency at that level. For example, at a given encoder/decoder level, $w^{1}$ denotes weights of the first transformer block, we can define the output $\mathbf{F}_{out}$ of that encoder/decoder level, given input features $\mathbf{F}_{in}$ as,
@@
We gradually increase the reuse frequency of the transformer block from top level to the bottom level, thereby increasing the network depth. The multi-level encoder systematically reduces spatial resolution while increasing channel capacity, enabling the extraction of low-resolution latent features from a high-resolution input image. Subsequently, the multi-level decoder gradually restores the high-resolution output from these low-resolution latent features. To enhance the decoding process, we integrate prompt blocks from PromptIR (Potlapalli et al., 2023), which learn degradation-aware prompts implicitly through the prompt generation module (PGM) followed by the prompt-interaction module (PIM). Unlike the existing PromptIR, our approach incorporates degradation-aware implicit prompt blocks at skip connections. This implicit prompting via skip connections enables the transfer of refined encoder features to the decoder side, thereby enhancing the restoration process. This fundamental correction in the network design gives us significant improvement for all-in-one image restoration tasks, as detailed in the experimental section (Sec. 5).
@@
At each encoder-decoder level, by adjusting the module weights reuse frequency ($f$), we obtain DyNet’s bulky and lightweight variants. Fig. 2 shows the core architecture of our DyNet. At each encoder-decoder level, we initialize weights ($w^{1}$) for the first transformer block and reuse them for the subsequent blocks. We vary the reuse frequency at each encoder-decoder level and obtain our bulkier and lightweight variants. DyNet comprises four encoder-decoder levels, with the larger variant (DyNet-L) employing reuse frequencies of $f=[4,6,6,8]$ at encoder-decoder: level 1 to level 4, respectively. In contrast, the smaller variant (DyNet-S) applies weights reuse frequencies $f=[2,3,3,4]$ for encoder-decoder level 1 to level 4, respectively. Further, we incorporate a prompt block at each skip connection, which enhances the encoder features being transferred to the decoder from the encoder. Overall, there are three prompt blocks: deployed one at each skip connection, each consisting of five prompt components each.
Figure 3: The proposed dynamic pre-training strategy is shown. Given a clean image, we create a degraded version by injecting noise (Gaussian or Random), JPEG artifacts and random masking within the same image. Two variants (small and large) of our DyNet are then trained concurrently to reconstruct the clean image from masked degraded inputs. Notably, the weights are shared between both variants since they are based on the same architecture but with an intra-network weight-sharing scheme with varying frequencies of block repetition. One of the two parallel branches is randomly activated in a single forward pass of the model (Flag being the binary indicator variable to show branch activation). The dotted lines show the network path is inactivated (Flag=0). An L1 loss is used to calculate pixel differences between reconstructed outputs and targets to update the shared weights of both branches. Both the inter and intra-model weight-sharing and random activation of branches lead to a significant reduction in GPU hours required for pre-training.